### v1

# R-Bot: An LLM-based Query Rewrite System

Abstract.
Query rewrite is essential for optimizing SQL queries to improve their execution efficiency without changing their results. Traditionally, this task has been tackled through heuristic and learning-based methods, each with its limitations in terms of inferior quality and low robustness. Recent advancements in LLMs offer a new paradigm by leveraging their superior natural language and code comprehension abilities. Despite their potential, directly applying LLMs like GPT-4 has faced challenges due to problems such as hallucinations, where the model might generate inaccurate or irrelevant results. To address this, we propose R-Bot, an LLM-based query rewrite system with a systematic approach. We first design a multi-source rewrite evidence preparation pipeline to generate query rewrite evidences for guiding LLMs to avoid hallucinations. We then propose a hybrid structure-semantics retrieval method that combines structural and semantic analysis to retrieve the most relevant rewrite evidences for effectively answering an online query. We next propose a step-by-step LLM rewrite method that iteratively leverages the retrieved evidences to select and arrange rewrite rules with self-reflection. We conduct comprehensive experiments on widely used benchmarks, and demonstrate the superior performance of our system, R-Bot, surpassing state-of-the-art query rewrite methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Query rewrite is designed to transform an SQL query into a logically equivalent version that is more efficient to execute, playing a crucial role in enhancing query performance in numerous practical scenarios. Despite its significance, the process of query rewrite is NP-hard (Zhou et al., 2021, 2023), meaning there is a vast collection of possible rewrite rules, and the number of potential rule combinations increases exponentially. This complexity makes identifying an effective combination of rules a challenging and laborious task. There are two main paradigms for addressing this challenge.
Heuristic-Based Methods. Some heuristic-based methods apply the rules in a fixed order derived from practical experience (e.g., PostgreSQL (pos, 2024)). However, they may not achieve optimal results for queries requiring different rule orders, thereby risking the omission of essential rewrite sequences. Furthermore, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to comprehensively explore various rule orders through heuristic acceleration. Nevertheless, they might overlook dependencies among rules, such as an initially inapplicable rule could be activated by another, leading to the potential neglect of vital rewrite sequences. Consequently, heuristic-based approaches are often criticized for their inferior quality.
Learning-Based Methods. To optimize query rewrite, learning-based methods have been proposed (Zhou et al., 2021, 2023). These methods employ neural networks that are trained on historical query rewrites to identify and apply the most advantageous rules for rewriting a query. However, learning-based approaches face criticism for their low robustness. For example, models trained through these methods struggle to adapt to unseen database schemas without undergoing additional training on hundreds of new query rewrite examples, which may not be readily available in many scenarios.
Recent advances in large language models (LLMs) have shown superiority in understanding natural language and code, as well as reasoning ability (Hu et al., 2024; Achiam et al., 2023; Du et al., 2024; Imani et al., 2023). As LLMs can capture the query rewrite capabilities by pre-training from database forums and codes, encompassing both the direct rewriting ability of applying a rule and the indirect rewriting ability to draw inspiration from multiple dependent rules, we can leverage LLMs to guide the query rewrite.
Table 1. R-Bot vs Existing Query Rewrite Methods.
Method Heuristic Fixed Order Heuristic Exploring Traditional Learning R-Bot LLM&Evidence
Train No No Yes No
#Rules High Low High High
Quality Low Low High High
Robust High High Low High
To realize this ambition, we aim to develop an LLM-based query rewrite system with two main advantages: $(i)$ High Quality. On one hand, our system can figure out the potential rules indirectly related to the query (e.g., ones that become applicable only after applying other rules). On the other hand, our system can understand the interrelations among rules and generate an effective sequence of rules for holistic improvements. $(ii)$ High Robustness. Compared to conventional learning-based approaches, our system using LLMs benefits from extensive pre-training, endowing it with superior adaptability and robustness.
However, directly utilizing LLMs for query rewrite proves to be ineffective due to their tendency for hallucination (Huang et al., 2023). For example, despite being pre-trained on a vast corpus of query rewrite data (e.g., Stack Overflow (sta, 2024)), employing the advanced LLM GPT-4 to directly rewrite queries in DSB benchmark (Ding et al., 2021) yielded only a 5.3% success rate, which is significantly low. This highlights two primary challenges associated with leveraging LLM for query rewrite.
C1: How to mitigate LLM’s factuality hallucination in query rewrite? It’s common for LLM to encounter confusion during query rewrite, suggesting an intuitive approach of guiding LLM with specific rewrite evidences (e.g., database Q&As, database manuals and codes, forum, etc) closely related to the query. However, several challenges arise from this approach. Firstly, LLM often struggles to interpret the raw rewrite evidence due to difficulties in aggregating fragmented knowledge across various rewrite documents and understanding complex query rewrite codes. Secondly, it’s crucial to sift through and identify the most beneficial rewrite evidences to serve as references for LLM, thereby steering it towards a more efficient rule selection for query rewrite.
C2: How to mitigate LLM’s faithfulness hallucination in query rewrite? LLM also encounters challenges in accurately analyzing complex queries, such as those containing multiple sub-queries, and in fully leveraging detailed rewrite evidences. Consequently, it becomes essential to develop a multi-step LLM rewrite algorithm that breaks down the query rewrite process into more manageable segments, thereby aligning with LLM’s capabilities for better performance.
Table 2. Example Rewrite Rules.
Rule Condition Transformation Matching Function
$r_{1}$ FILTER_SUB_QUERY_TO_JOIN
Scalar, “IN”, or “EXISTS” sub-
query in “WHERE” clause.
Transformed to join on
the correlated column.
b -¿ b.operand(Filter.class)
.predicate(containsSubQuery);
$r_{2}$ FILTER_INTO_JOIN
Filter condition with column
from one join side.
Push down condition to
filter non-nullable side.
b -¿ b.operand(Filter.class)
.oneInput(Join.class);
$r_{3}$ AGGREGATE_PULL_UP_CONSTANTS
Some “GROUP BY” key is
constant across rows.
Remove constant key.
Constant may project.
b -¿ b.operand(Aggregate.class)
.predicate(hasConstantExps);
To tackle the above challenges, we propose R-Bot, an LLM-based query rewrite system designed with a systematic approach. First, we gather and prepare rewrite evidences from diverse sources, including integrated rewrite rules aggregating from rewrite documents and summarizing from the complex rule codes, as well as high-quality Q&As from database forums (addressing C1). Second, for an input SQL query, we propose a hybrid structure-semantics method for retrieving pertinent evidences, including rewrite rules by matching functions and rewrite Q&As with both query structure and rewrite semantics similarities (addressing C1). To enable LLM to comprehend the rewrite evidences, we synthesize this information into rewrite recipes, which detail in natural language how to utilize the Q&As and rewrite rules for rewriting an SQL query, and we retrieve most relevant rewrite recipes to guide LLMs for query rewrite (addressing C2). Third, we design a step-by-step LLM rewrite algorithm, which guides LLM to iteratively utilize the rewrite recipes to refine its rewrite rule selection and ordering, possibly choosing a more promising rule sequence by reflecting and self-improving the rewrite process (addressing C2).
Contributions. In summary, we make the following contributions.
$(1)$ We develop an LLM-based query rewrite system, R-Bot, which can select an effective rewrite rule sequence to guide query rewrite engines to rewrite a query (see Section  3).
$(2)$ We design a multi-source rewrite evidence preparation pipeline, including clustering-based document re-organization and hierarchical code summarization for rewrite codes (see Section 4).
$(3)$ We propose a hybrid structure-semantics retrieval method for retrieving relevant rewrite evidences (see Section 5).
$(4)$ We propose a step-by-step LLM rewrite method that iteratively leverages the retrieved Q&As and rewrite recipes to select and arrange rewrite rules with self-reflection (see Section 6).
$(5)$ Our experimental results on widely used benchmarks demonstrate that R-Bot can significantly outperform existing state-of-the-art query rewrite methods (see Section 7).
2. Preliminaries
In this section, we first formalize the problem of query rewrite (see Section 2.1). Then we introduce fundamental LLM concepts used in R-Bot (see Section 2.2).
2.1. Query Rewrite
Rewriting a query involves numerous query transformations, and we typically identify and summarize common transformation patterns into rewrite rules. For instance, a query rewrite rule could involve replacing an outer join with an inner join when they are equivalent. The composition of these rules offers the flexibility to accommodate a wide range of query rewrite requirements.
Definition 2.1 (Rewrite Rule).
A rewrite rule $r$ is denoted as a triplet $(c,t,f)$, where $c$ is the condition to use the rule, $t$ is the transformation to be applied, and $f$ is a matching function used for evaluation. If a query $q$ satisfies the condition $c$ as determined by the matching function $f$, then the transformation $t$ can be applied to the query $q$, resulting an equivalent rewritten query $q^{[r]}$.
For example, Table 2 shows some query rewrite rules. Since the column “comm” with condition “comm=100” in the SQL query satisfies the condition “some ‘GROUP BY’ key is constant across rows” of rule $r_{3}$, we can apply the rule and remove the column “comm” from the “GROUP BY” clause.
Numerous rewrite rules have been pragmatically incorporated into database products, as evidenced in existing literature and products such as PostgreSQL (pos, 2024), MySQL (mys, 2024), Apache Calcite (Begoli et al., 2018). However, when applying the rules to rewrite a query, it is often cumbersome to decide the best rule sequence for two primary reasons. First, since a rewrite rule sometimes degrades the query (e.g., $q^{[r]}$ with higher execution latency than $q$), we should examine whether or not to use the rule. Second, it is also important to decide the order of applying the rules. For instance, applying one rule may render another rule obsolete. Thus query rewrite aims to find an optimal rule sequence to rewrite a query in order to minimize the execution cost of the rewritten query, which is formulated as below.
Definition 2.2 (Rule-based Query Rewrite).
Consider a query $q$ and a set of rewrite rules $R$. Assume $\alpha$ is a sequence of certain rewrite rules selected from $R$. Let $q^{\alpha}$ denote the rewritten query by sequentially applying the rules in $\alpha$ to rewrite the query $q$. The objective of query rewrite is to obtain an optimal rule sequence $\alpha^{*}$, such that the execution cost of $q^{\alpha^{*}}$ is minimized among all possible rewritten queries of $q$.
The query rewrite problem has been proven to be NP-hard (Zhou et al., 2021). Traditional methods cannot select high-quality rules. To address this limitation, we advocate for utilizing LLM to select rewrite rules. To address the hallucination problem of LLMs (Huang et al., 2023), we perform an offline stage to extract query rewrite evidences, including rewrite Q&As and rewrite rule specifications, and store them as Q&A repository and rewrite rule specification repository respectively. During the online phase, given an input SQL query, we retrieve relevant Q&As and rule specifications, and generate rewrite recipes, which outline how to rewrite a query using rewrite Q&As and rule specifications, thereby assisting LLMs in comprehending the rewrite evidences. With the assistance of the rewrite recipe, LLMs are then guided through a step-by-step process to judiciously select and apply rewrite rules to the query. Next we formally define these notations.
Definition 2.3 (Rewrite Q&A).
A rewrite Q&A includes a query rewrite question and a rewrite answer on how to rewrite the query in natural language.
Definition 2.4 (Rewrite Rule Specification).
A rewrite rule specification is used to describe a rewrite rule $(c,t,f)$ using natural language, which is also a triplet $(nc,nt,f)$, where $nc$ describes the condition and $nt$ describes how to apply the transformation in natural language.
For example, Figure 1 shows a rewrite rule specification derived from Calcite code of rule “AGGREGATE_PULL_UP_CONSTANTS”, which explains the condition and transformation in natural language. Besides, the rule specification will be retrieved as evidence in this example, and we can use this evidence to rewrite the query.
Based on the general concepts, we further introduce definitions related to particular query rewrite, including rewrite Q&A recipe and rewrite rule specification recipe.
Definition 2.5 (Rewrite Q&A Recipe).
Given a query $q$ and a Q&A, a Q&A recipe provides instructions in natural language on how to utilize the Q&A to rewrite the query $q$.
Definition 2.6 (Rewrite Rule Specification Recipe).
Given a query $q$ and a rewrite rule specification, a rewrite rule specification recipe describes how to use the rewrite rule specification to rewrite the query $q$ in natural language.
For instance, Figure 1 shows examples of rule specification recipe and Q&A recipe.
Figure 1. Query Rewrite Example: the sequence of rewrite rules $[r_{1},r_{2},r_{3}]$ is found by R-Bot based on retrieved rewrite rule specifications and rewrite Q&As.
In this paper, we focus on how to prepare the Q&A repository and rewrite rule specification repository (see Section 4), how to generate Q&A recipe and rule specification recipe (see Section 5), and how to use them to rewrite a query (see Section 6).
2.2. Large Language Models
We introduce some fundamental concepts of LLMs, including LLM prompting and retrieval augmented generation (RAG), which are pivotal in composing an evidence-driven query rewrite system.
LLM Prompting. After instruction tuning on large corpus of human interaction scenarios, LLM is capable of following human instructions to complete different tasks (Ouyang et al., 2022; Dong et al., 2022). Specifically, given the task input $X$, we concatenate a prompt $P$, denoted as $X^{\prime}=P\bigoplus X$, which represents a comprehensive task description. LLM then fulfills the task, and generates the task output. Note that pre-training on trillions of bytes of data enables LLM to generalize to diverse tasks by simply adjusting the prompt  (Achiam et al., 2023; Khattab et al., 2023).
Retrieval Augmented Generation. During completing tasks with prompting, LLMs often generate erroneous or meaningless responses, i.e., the hallucination problem (Huang et al., 2023). To mitigate the problem, retrieval augmented generation (RAG) is utilized as an advanced LLM prompting technique, which is composed of three steps (Gao et al., 2023): $(i)$ indexing, which prepares chunks of the task-specific knowledge source with embedding vector index; $(ii)$ retrieval, given a user query, using the index to retrieve the most relevant chunks to the query; $(iii)$ generation, guiding LLM to generate answers with the retrieved chunks as a prompt.
Figure 2. Overview of R-Bot.
However, existing RAG and prompting techniques face limitations when applied directly to query rewrite for two primary reasons. First, there is absence of embedding methods capable of accurately evaluating the similarity between Q&As and the input query. For example, the conventional RAG technique relies on text embeddings for both the Q&A and the query, capturing only their semantic similarity while neglecting crucial structural information pertinent to query rewrite. To address this gap, we propose a hybrid structure-semantics approach for retrieving relevant Q&As, as detailed in Section 5. Second, the multitude of rules presents a significant challenge for LLMs, as directly arranging these rules can lead to serious hallucination problems. This necessitates the development of a task-specific, step-by-step LLM algorithm that breaks down the query rewrite process into simpler, more manageable stages, as elaborated in Section 6.
3. The Overview of R-Bot
Our R-Bot system includes an offline stage and an online stage, as described in Figure 2.
Offline Rewrite Evidence Preparation. This stage aims to extract rewrite Q&As from the Web and rewrite rule specifications from rewrite codes and database documents, subsequently storing them in Q&A repository and rule specification repository, respectively. These resources will be utilized to assist in rewriting an online SQL query. Note that rewrite rule specifications are more general that can be applied to multiple SQL queries. In contrast, rewrite Q&As are usually specific and are applied to a particular SQL query.
We extract rewrite rule specifications from two types of resources. $(i)$ Database menus and documents. Given that query rewrite evidence is often dispersed across various documents, sections, and paragraphs, it becomes necessary to aggregate these evidences from diverse sources through semantic clustering and distill them into a coherent rewrite rule; $(ii)$ Rule codes. Considering the complexity of the code, characterized by its intricate nested calls, we employ a hierarchical strategy to streamline the code from simple to complex. This approach involves initially summarizing straightforward functions, followed by recursively summarizing more complex functions. In this process, symbol references are clarified using functions that have already been summarized. Ultimately, this method results in a concise summary of the overall rule code, effectively condensing the rule into a simplified format.
We extract Q&As from website Q&As (e.g., Stack Overflow (sta, 2024)). Given the wide range of topics and variable quality of these website Q&As, we filter out high-quality Q&As related to query rewrite, by SQL selections (e.g., on question tags) and LLM filtering.
Additionally, we construct efficient indexes to enhance the performance of Q&A and rule specification retrieval. We will discuss the detailed techniques in Section 4.
Online LLM-Guided Query Rewrite. Given an online SQL query, this stage retrieves pertinent Q&As and rule specifications, leverages them to rewrite the query, and offers reflections that not only complete the rewrite process but facilitate its further refinement.
(1) Hybrid Structure-Semantic Evidence Retrieval. For an online SQL query, we retrieve relevant evidences from rewrite rule specification repository and Q&A repository.
  1. (i)
    Rule specification retrieval. As the rule specification has matching conditions and matching functions, we can easily retrieve the relevant rule specifications whose matching functions are satisfied by the input SQL query;
  2. (ii)
    Q&A retrieval. There are two types of Q&As that could potentially be relevant to the SQL query. First, the SQL questioned in Q&A structurally matches the input SQL query. We require a structure-aware matching method to retrieve such Q&As. To this end, we propose a query structure embedding composed of $(i)$ the query template embedded by pre-trained embedding, and $(ii)$ one-hot embedding which represents the query’s matched rule specifications to assess the structural similarities. Second, Q&A semantically matches the input SQL query, such as the rewrite explanations provided in the answer section of the Q&A. However, SQL usually has no natural language rewrite explanations. To this end, we can leverage the retrieved rule specification to identify relevant Q&As based on their semantics. We propose a semantics embedding method to embed the rule specification and Q&A, and then assess their similarities based on a similarity function (e.g., $L^{2}$-distance (chr, 2024)). Lastly, to retrieve Q&As relevant to the input query both structurally and semantically, we merge the query’s structural embedding with its semantic embedding into a unified representation. We then build an embedding index for Q&A repository offline. For an online query, we utilize this index to efficiently retrieve top-$k$ relevant Q&As.
For retrieved rule specification and Q&A of the input SQL query, we leverage LLM to generate SQL-aware rule specification recipe and Q&A recipe that describe how to utilize them to rewrite this SQL query. We will discuss the technical details in Section 5.
The retrieved rule specification recipe and Q&A recipe will be used to reformulate the SQL query in the following steps.
(2) Step-by-Step LLM Rewrite. Given the rewrite rules supported by the query rewrite engine (e.g., Apache Calcite (cal, 2024)), we direct LLM to select pertinent rules to rewrite. As there are many rules, if we directly instruct LLM to arrange a rule sequence, LLM can encounter serious hallucination problem. To mitigate this problem, we propose a step-by-step LLM rewrite method that decomposes rule-based query rewrite into several simpler steps.
  1. (i)
    Q&A-based rule selection and ordering. Given the rewrite rules and a sorted sequence of Q&As retrieved from the previous step, ranked by the relevance score, we select and rank the rules. We first initialize the score of each rule as 0. Then for each pair of a rule and a Q&A, we use LLM to evaluate whether they are relevant, i.e., whether the rule is applicable in light of the Q&A. If applicable, we increase the score of the rule by the score of this Q&A. Then by enumerating all the pairs of rules and Q&As, we can get the final score of each rule and rank the rules based on the final score. Since the relevance can be evaluated by LLM offline, this step can be efficiently executed by algorithms.
  2. (ii)
    Recipe-based rule selection and ordering. Building on the preliminary rule sequence established in the previous step, we utilize LLM to sift through and exclude any rules that do not align with the recipe. A straightforward way is to enumerate each pair of recipes returned by step (1) and rules returned by step $(2.i)$, and ask LLM to evaluate their relevance and rank the rules. However, there are two limitations. First, it may overlook the rule relevance, since a recipe may encompass multiple rules. Second, since the recipe is generated from queries and cannot be evaluated offline, assessing each pair with LLM becomes costly. To address this issue, we propose an effective filtering method to efficiently select the rules.
  3. (iii)
    Rule-based rewrite. Based on the selected rules, we input them to the query rewrite engine to rewrite the query.
We will discuss the technical details in Section 6.
(3) Rewrite Reflection. It provides rewrite reflections to either further refine the query or to finalize the rewrite process. It has two reflection resources. The first involves getting the cost of the rewritten query from databases, comparing it with the cost of the query prior to its rewrite, and returning complete if the cost of the rewritten query is smaller; continue otherwise. The second involves asking LLMs to check whether or not all the rewrite recipes are realized by the query rewrite in this step, and returning complete if yes; continue otherwise. So there are four possible reflections.
  1. (i)
    complete, complete: It finalizes the rewrite process and returns the rewritten query.
  2. (ii)
    complete, continue: It further rewrites the query by jumping to step (1) with the previously rewritten query as the input. Specifically, it starts a new round of LLM-guided query rewrite, where new rewrite evidences can be retrieved and new rules can be selected based on the previously rewritten query.
  3. (iii)
    continue, complete: It further refines the query by jumping to step $(2.ii)$, focusing on reordering the existing set of rules. Specifically, besides the recipes and the rules, we further input the rules actually used in the previous rewrite process, and instruct LLM to prioritize the unused rules.
  4. (iv)
    continue, continue: This approach integrates elements from branches $(3.ii)$ and $(3.iii)$. Initially, it defaults to branch $(3.iii)$ unless this path is revisited excessively, surpassing a predefined threshold. Under these circumstances, given that branch $(3.iii)$ has exhaustively explored the rules for the current query, the process transitions to branch $(3.ii)$. This shift initiates a new cycle, aiming to further refine the query.
4. Rewrite Evidence Preparation
We discuss how to extract and standardize rewrite evidences from diverse rewrite sources. This evidence is crucial for crafting a comprehensive rewrite recipe that guides the LLM rewrite process. We explain respectively how to prepare rewrite rule specifications (see Section 4.1) and rewrite Q&As (see Section 4.2).
4.1. Rewrite Rule Specification Preparation
The rewrite rule specification clearly outlines, in natural language, the condition for use, the query transformation operations to be executed, and the matching function used for evaluation. It contains three key components. $(i)$ “condition”: a prerequisite that a query must fulfill to utilize the rule; $(ii)$ “transformation”: detailed steps for transforming the query into an equivalent form that is optimized for more efficient execution; and $(iii)$ “matching function”: an executable function that outputs ‘1’ if the input query matches the rule; and ‘0’ otherwise.
Generating rewrite rule specifications is challenging due to the considerable effort needed to distill and synthesize information from various sources into a concise format. This process includes summarizing extensive rewrite codes, which may span thousands of lines and feature complex structures, as discussed in Section 4.1.1. Additionally, it integrates crucial but scattered information from documents on rewrite rules, as outlined in Section 4.1.2.
Figure 3. Transforming Rule Code into Rule Specification.
4.1.1. Transforming Rule Code into Rule Specification
Given that some query rewrite engines, such as Apache Calcite (cal, 2024), are not accompanied by comprehensive documentation, we are compelled to decipher the rewrite rules directly from the raw code. This process involves navigating through complex code structures that include intricate nested calls. To address this challenge, we introduce a hierarchical rule code summarization method, as illustrated in Figure 3. Our approach begins with the construction of a rule code structure tree, emanating from the rule’s main function. In this tree, each node represents a symbol declaration (e.g., functions, variables, classes), while each edge denotes a symbol reference relationship. Progressing through the structure, we methodically summarize the declaration code, moving from simple to more complex elements and clarifying symbol references using previously summarized symbols. With the comprehensive summary of the root node at our disposal, we guide LLM to convert this information into a standardized rewrite rule specification.
Rule Code Structure Analysis. Given the symbol references of the rule code, it is crucial to clarify the symbol declarations before summarizing the rule code. We first build a root node representing the main function of the rule. Then, we use code analysis tools (e.g., JavaSymbolSolver (jav, 2018)) to associate the symbols in the main function with their corresponding declarations, which are children nodes of the root node. If the declaration of some nodes also accesses other unseen symbols, we further resolve its symbol references. We recursively expand the nodes until reaching built-in symbols. In this way, we obtain a rule code structure tree, where each node represents a symbol declaration and each edge represents a symbol reference relationship.
Hierarchical Rule Code Summarization. The complex rule code structure poses two challenges for LLM code summarization. First, given the necessity of every declaration in the rule code structure, if we directly input them to LLM, the long context can greatly degrade LLM performance (Liu et al., 2024a; Li et al., 2023). Second, the considerable width and depth of the structure (e.g., tens) further increase the reasoning complexity for LLM. We thus propose a hierarchical rule code summarization method. First, if a node declaration already has detailed comments, there is no need to summarize the node. Second, for the leaf node without comments, it can be directly summarized by LLM due to the simple code structure and absence of unfamiliar symbols. Third, for the non-leaf node whose children are already summarized, the symbol references in the declaration code can be clarified by its children summaries. Specifically, we insert the symbol summaries as comments into the declaration code, which enables LLM to accurately summarize the code. We repeat the process until we obtain the summary of the root, which represents the whole rule code.
Rule Specification Regularization. To facilitate LLM understanding, we regularize the rule code as a standard rewrite rule specification. Specifically, we use LLM to extract the condition and transformation of the rule, using the prompt $p_{reg}=$“Given a rewrite rule code summary, your task is to extract the rewrite rule that explains completely and detailedly the condition and transformation.”.
4.1.2. Transforming Rewrite Document into Rule Specification
Considering the variety of rewrite documents, such as those for PostgreSQL and MySQL, note that sections within a single document may cover rewrite rules that bear weak relation to one another. For example, optimizations for the “WHERE” clause might discuss both constant folding and index utilization without clear interrelation. Additionally, components complementary to a rule can be dispersed across different documents. For instance, while a MySQL document might detail conditions conducive to acceleration via index utilization, a separate PostgreSQL document could highlight how certain column transformations might inhibit the use of indexes. Together, these insights from disparate sources can contribute to forming a comprehensive rewrite rule specification.
To address this, we propose a clustering-based document re-organization method. First, we use LLM to extract rewrite rules from rewrite documents. Second, for extracted rules, we cluster the correlated ones together into one group, where we evaluate their semantics similarities by their text embeddings (e.g., SBERT (Reimers and Gurevych, 2019)). Third, we use LLM to summarize each rule cluster, and transform each cluster summary as a regularized rewrite rule specification.
Rule Extraction. We use LLM to extract rewrite rules from the rewrite documents in two steps. First, if we directly input all the documents to LLM, it often overlooks important details in the middle of the extremely long context (e.g., 100k) (Liu et al., 2024a; Li et al., 2023). We thus split the documents into structured blocks (e.g., sections, sub-sections) that each can be effectively processed by LLM. Second, we instruct LLM to extract rewrite rules from the split blocks. To mitigate the hallucination problem (Rawte et al., 2023), we require LLM to locate supporting content in the source document. If no such content can be found, we can deem the extraction low quality and repeat LLM extraction.
Rule Clustering. To identify pertinent rules among different chapters and documents (e.g., condition push-down involved in both PostgreSQL and MySQL documents), we adopt a clustering algorithm based on Gaussian Mixture Models (GMMs) (Sarthi et al., 2024). Specifically, the semantics of each rule is represented as a embedding vector (e.g., using multi-qa-mpnet-base-cos-v1 (Reimers and Gurevych, 2019; mul, 2024)). It then fits the vector distribution with a weighted combination of $M$ Gaussian distributions. Each distribution forms a cluster, assigning rules to the cluster with the highest density of their embeddings. To address the limitations of GMMs in high-dimensional spaces, it adopts Uniform Manifold Approximation and Projection, which reduces dimensionality by approximating the local manifold of the data (McInnes et al., 2018). Besides, to determine the optimal cluster number $M$, it adopts Bayesian Information Criterion, which aims to maximize the overall distribution likelihood $\hat{L}$ while minimizing the cluster number (Schwarz, 1978). This is achieved by minimizing $ln(N)m-2ln(\hat{L})$, where $N$ is the number of rules, and $m$ is proportional to the cluster number $M$.
Rule Specification Regularization. We use LLM to obtain standardized rule specifications from the rule clusters. First, we summarize each cluster using LLM with a prompt $p_{summ}=$“Given the rewrite rule components, your task is to summarize them into one paragraph, and your summary should include as many details as possible.” Then, we further use LLM to transform the summary to a regularized rewrite rule specification, by extracting the conditions and transformations of the rule.
4.2. Rewrite Q&A Preparation
The rewrite Q&A showcases query rewrite of how to transform a particular SQL query to an optimized one, which is composed of two parts: $(i)$ “question” that denotes a query rewrite request; $(ii)$ “answer” that provides the query transformations for the question. There are plenty of rewrite Q&As within online database community forums (e.g., millions at Stack Overflow (sta, 2024)), which may cover more flexible and practical cases beyond the rewrite rules.
Since the website Q&As are mixed of diverse topics and qualities, we use a hybrid method to filter out the high-quality Q&As on query rewrite. First, we select Q&As by their question tags (e.g., “query-optimization”) and community comments (e.g., a Stack Overflow score higher than 3). Second, we use LLM to examine the topics of these Q&As, and select those related to query rewrite.
5. Hybrid Structure-Semantics Evidence Retrieval
Considering the existence of vast repositories of rule specifications and Q&As, only a minimal fraction of these resources is pertinent to an online SQL query. Thus, there is a need to identify the relevant ones both effectively and efficiently. In this section, we introduce how to retrieve relevant rewrite evidences (including rule specifications and Q&As). First, we retrieve relevant rule specifications using a function-based rule retrieval method (see Section  5.1). Second, we retrieve relevant rewrite Q&As using a hybrid structure-semantics method, with both query structures and rewrite semantics aligned with the input query (see Section  5.2). Lastly, we generate tailored rewrite recipes for the input query by leveraging both the retrieved rule specifications and Q&As (see Section  5.3).
5.1. Rewrite Rule Specification Retrieval
Given an input SQL query $q$, we examine each rule specification and apply its associated matching function to get a boolean value which indicates whether the condition of the rule is satisfied. If the result is true, the corresponding rule specification $rs$ is retrieved. Then, we use LLM to generate a rule specification recipe, using a prompt $p_{rule\_spec\_recipe}=$“Given an SQL query ‘$q$’ and a rewrite rule specification ‘$rs$’, your task is to explain concisely and detailedly how the rule applies to the query, by specifying (1) the SQL segments matched by the condition, and (2) the transformation of the rule.”
Figure 4. Structure-Semantics Embeddings in Q&A Retrieval.
5.2. Rewrite Q&A Retrieval
There are two types of Q&As that could potentially enhance the query rewrite. Firstly, the questions within Q&As exhibit a high structural similarity to the input query, indicating that the answers of Q&As have the potential to assist in rewriting the query. Second, the answers within Q&As demonstrate a high semantic similarity to the input query. This suggests that the Q&As are capable of addressing similar issues or bottlenecks present in the input query, such as those involving a sub-query with an aggregate function.
Structure-Semantics Embeddings. To effectively identify these two types of Q&As, we propose structure-semantics embeddings for both SQL queries and Q&As (see Figure 4). (1) SQL Query Embedding. First, we introduce a structure-aware query embedding strategy, including $(i)$ generating query templates embedded by pre-trained embedding (e.g., text-embedding-3-small (ope, 2024)), and $(ii)$ generating a one-hot embedding to indicate which rule specifications match the SQL query. The two embeddings capture the essential structural features to query rewrite (see Section 5.2.1). Second, we propose a semantic matching method designed to discern the semantic similarity between the SQL query and Q&As. However SQL may not encapsulate sufficient semantic information. Fortunately, the retrieved rule specification recipe of SQL in Section 5.1 encompasses these semantics, enabling us to derive an embedding from it. Third, we concatenate the structural and semantic embeddings to obtain a combined structure-semantics embedding. Given the possibility of retrieving multiple rule specification recipes, each SQL query has the potential to generate multiple embeddings. (2) Rewrite Q&A Embedding. Similarly, we generate embeddings of Q&As. For each Q&A, its embedding includes a structure-aware embedding of its query in the question part and a semantic embedding of its answer part. These embeddings allow us to identify Q&As that exhibit a high similarity to the input query in terms of their embeddings.
Structure-Semantics Q&A Retrieval. To optimize performance for structure-semantics Q&A retrieval, a unified structure-semantics embedding index (e.g., HNWS (chr, 2024)) is constructed for Q&As offline. Given an SQL query, we generate its structure-semantics embedding and retrieve relevant Q&As using the index (see Section 5.2.2).
5.2.1. Structure-Aware SQL Query Embedding
Existing query embedding models (Tang et al., 2022; Zhao et al., 2022, 2023; ope, 2024) cannot be directly applied for embedding the query structure for query rewrite due to limitations in two key aspects. On one hand, the query-specific embeddings mostly adopt small-scale neural networks (e.g., lower than 0.1B) fine-tuned on a specific dataset (Zhao et al., 2022, 2023) or a specific workload (Tang et al., 2022), and thus struggle to generalize to arbitrary scenarios. On the other hand, the general text embeddings (e.g., text-embedding-3-small (ope, 2024)) cannot be directly used for query structure embedding, which are sensitive to irrelevant query information for three reasons. $(i)$ The identifiers (e.g., table/column names) and literals in the query often disrupt characterization of the query structure. For example, while renaming the schema in a query can maintain the integrity of the query rewrites, this action results in a version with significantly altered semantics and it generates an embedding distinctly different from that of the original query; $(ii)$ Since query rewrites typically focus on transforming only certain parts of the query (e.g., a sub-query), disregarding other irrelevant parts can help clarify the query structure; $(iii)$ The text embedding models cannot guarantee commutative invariance, implying they may generate dissimilar embeddings for expressions such as “$p_{1}\ and\ p_{2}$” and “$p_{2}\ and\ p_{1}$”. These can be addressed through carefully designed textual transformations.
To address the aforementioned challenges, we propose a structure-aware query embedding method, where each SQL query embedding is composed of two parts. The first part distills the core query structure as query templates, and uses pre-trained text embedding to embed them. The second part employs a one-hot encoding strategy to indicate whether the query matches a rule specification, with ‘1’ representing a match and ‘0’ indicating no match. The width of this embedding corresponds to the total number of rule specifications. When comparing the two query structure embeddings, the first one captures a more global representation of the query structure, whereas the second focuses on more local expressions.
Dataflow Based Query Template Embedding. To effectively capture the essential SQL features for query rewrite, we initially reformulate the SQL query into composite dataflows that detail operations on the involved identifiers. Specifically, given the structure and semantics of a query $q$, a dataflow $\delta$ is a list of SQL operations sequentially performed on one or multiple tables or columns in the query, which corresponds to an SQL segment of the query, e.g., logical expression, mathematical expression, “WHERE” clause, “GROUP BY” clause, “FROM” clause, “JOIN” clause, “SELECT” list, sub-query, etc. For instance, we show sample dataflows for the input query in Figure 1, concentrating on the column “$bonus.sal$”:
$\{sal,SELECT\ MAX(sal),emp.sal=ANY\ (SELECT\ MAX(sal)\ ...)\}$
where “$sal$” is first input from “$bonus$” table, projected with an aggregate function “$MAX(\cdot)$”, and then compared with the column “$emp.sal$” from the outer query.
Since we focus on logical query rewrite, dataflows can be considered independent if they do not involve common identifiers. Then, if we examine the rewrite rule from dataflow perspective, the condition must restrict the pattern of certain dataflows associated with a specific identifier. Otherwise, there are scarcely any potential query rewrites within a set of independent dataflows. Consequently, we can generate query templates that concentrate on the SQL operations related to each particular identifier.
To ensure that the generated query template remains unaffected by schema renaming, we propose an isolated masking method. If a query template involves multiple identifiers, swapping any two identifiers alters its semantics and disrupts the query template embedding. Thus, we derive multiple query templates from the query, with each template preserving only one identifier while masking the others. We then refine the query templates in three steps. First, since a complex query containing many identifiers can derive a large number of query templates, we opt to select only a small part of them as representative. Specifically, if the SQL operations associated with an identifier are potential to be matched by some rewrite rules, we select the corresponding query template as representative, since it concentrates on those potential SQL operations on the identifier. Second, we simplify the query template by eliminating the SQL operations that do not involve the key identifier. Third, for the commutative SQL operands, we sort them lexically, thereby rendering the query template resistant to potential commutations.
Step 1: Potential Identifier Selection. Considering the SQL operations associated with the identifier, we observe that if the identifier appears multiple times in some mutually exclusive dataflows, it can possibly trigger query rewrite by utilizing the correlations of the dataflow operations. For instance, the twice appearance of column “$a$” in expressions “$a<b$” and “$a=5$” can indicate constant folding, which replaces “$a<b$” with “$5<b$”. Then, given an SQL query, we propose two ways to count the frequency of identifier occurrences. $(i)$ Column appearances: By traversing the query, we count how often each column appears. Note that columns appearing as direct projections (e.g., “SELECT a”) are not counted, as this does not involve any meaningful SQL operations. $(ii)$ Table appearances: We count the appearances of the table where the table is utilized either in the “FROM” clause or within a “JOIN” operation. With the frequency of tables and columns, we focus on identifiers that appear more than once within the query. We then select their corresponding query templates as representative, ensuring we capture the most significant patterns for query rewrite.
Step 2: Query Template Reduction. Given the single identifier preserved in the query template, we standardize it as “$table$” for table and “$column$” for column. Moreover, to remove irrelevant query information, we mask other identifiers with “_” and the literals with “?”. We make a distinction between identifier mask and literal mask, since they indicate different query rewrite potentials. For instance, we can preserve the constant folding pattern in the query template with “$column<\_\ AND\ column=?$”, which means the constant equality can be transferred to other conditions involving “$column$”. Then, we further simplify the query template following three steps. $(i)$ If any dataflow in the query template involves only masked literals “?”, no identifiers, and no non-deterministic functions, the dataflow can be regarded as a literal and replaced with literal mask “?”. $(ii)$ If any dataflow involves some identifiers but no explicit identifiers, we replace it with identifier mask “_”, as it is independent to the potential identifier. $(iii)$ For the dataflow of clause like “$t_{1}$ JOIN $t_{2}$ ON $c$”, if the join condition “$c$” and a join table (e.g., “$t_{2}$”) are both masked with “_”, we replace the clause with “$t_{1}$” to remove irrelevant details. These steps are repeated to streamline the query template until it cannot be further matched, leaving a structure that retains only the essential core related to the potential identifier.
Step 3: Commutative Invariance Guarantee. We further transform the query template for commutative invariance. Specifically, we search for commutative operators in the query template (e.g., addition, set intersection), and sort their operands lexically (e.g., “$column<\_$” before “$column=?$”). Equipped with the refined query templates, we then obtain query template embeddings with pre-trained text embedding (e.g., text-embedding-3-small (ope, 2024)).
One-Hot Embedding for Matched Rule Specifications. Given an SQL query, its matched rule specifications also reflect the query structure. For instance, queries matching “SUB_QUERY_TO_JOIN” rule likely contain similar query structures surrounding particular sub-queries. As shown in Section 5.1, we first examine the rule specifications with their associated matching functions. Following this, we construct a one-hot encoding from the rule matching results, where each position corresponds to a rule specification. If a rule specification matches the query, its corresponding position is ‘1’; ‘0’ otherwise. The one-hot embedding for matched rule specifications can capture local structural information indicated by the rule condition, which is complementary to the global structures captured by query template embedding.
5.2.2. Structure-Semantics Q&A Retrieval
First, we build a hybrid structure-semantics embedding index for Q&A repository in three steps. $(i)$ We extract the queries in the Q&As, generate query templates, and embed them with pre-trained text embedding (e.g., text-embedding-3-small (ope, 2024)). We also match the queries with rule specifications to build one-hot embedding for matched rule specifications. $(ii)$ We refine the Q&A by preserving the semantics of query rewrite and eliminating irrelevant text with the help of LLM, then embedding the streamlined information using a pre-trained text embedding model (e.g., text-embedding-3-small (ope, 2024)). $(iii)$ We normalize the three embeddings as unit vectors, concatenate them to form a holistic embedding, and insert them into the index. Next, for an online SQL query, we similarly generate its embedding and identity the most relevant Q&As with top-$k$ embedding similarities.
Every SQL query may be associated with multiple templates, leading to multiple structural embeddings. Similarly, each SQL query encompasses multiple specification recipes, resulting in multiple semantic embeddings. We concatenate these embeddings for all possible combinations and, for each concatenated embedding, we utilize the index to identify top-$k$ Q&As with the highest similarities. To combine the retrieved results from multiple embeddings, we adopt the Reciprocal Rank Fusion (RRF) method (Cormack et al., 2009). Specifically, for each retrieved list corresponding to an embedding, we assign a score to the $i$-th Q&A in the ranked list as, $s(qa_{i})=\frac{1}{\alpha+i}$, where $\alpha$ is 60 by default. The RRF score of a Q&A across all the retrieved lists is calculated by summing its scores from each list (assigning a score of 0 if the Q&A is absent in a list). We then identify the Q&As with highest top-$k$ RRF scores as the final selection, which comprehensively captures the structure and semantics relevance.
5.3. Rewrite Recipe Generation
For an input SQL query $q$, since the retrieved Q&A $qa$ is used to rewrite similar but different queries, we first use LLM to generate rewrite recipes describing how to rewrite the input query inspired by the Q&A, using the prompt $p_{qa\_recipe}=$“Given an SQL query ‘$q$’ and a rewrite Q&A ‘$qa$’, your task is to propose some strategies on rewriting the query, by (1) transferring the Q&A strategy to the query, and (2) explaining the strategy detailedly.” Second, to integrate both rule specification recipes (see Section 5.1) and Q&A recipes, we condense those that are closely related and eliminate any duplicates. Thus, we utilize LLM to cluster them by their semantic similarity, and summarize each recipe cluster concisely into a single recipe.
6. Step-by-Step LLM Rewrite
Given a rule-based query rewrite engine (e.g., Apache Calcite (cal, 2024)), we utilize the retrieved Q&As and derived rewrite recipes to assist LLM in selecting appropriate rewrite rules from the engine to rewrite the input SQL query. Given that the number of possible rule sequences grows exponentially with the number of rules, LLMs are susceptible to errors when choosing from a vast array of rules. To address this challenge, we introduce a step-by-step filter-refinement method designed to meticulously select high-quality rules. Specifically, we start with a filtering step by employing an efficient Q&A-based method for rule selection and ordering, allowing us to preliminarily arrange the rules. Next, we instruct LLM to select the rules and arrange the order according to the recipes. Then, we feed the selected rules into the query rewrite engine to rewrite the query with the rules. Lastly, we evaluate the outcomes of the rewrite process to determine whether further refinement is needed or if the query rewrite can be considered complete.
Step 1: Q&A-based Rule Selection and Ordering. Given the rewrite rules and the input query, we filter the rules and arrange the order in three steps. First, we select the rules directly matched by the query, and assign them a relevance score according to their transformation to the query. Second, we select the rules indirectly relevant to the query using the retrieved Q&As. Based on the retrieval scores of the Q&As, we assign each indirectly relevant rule with a relevance score. Third, we rank the rules by their relevance scores in a descendent order, which serves as an initial order.
$(i)$ Since the rules have matching functions, we first filter the rules whose matching functions are satisfied by the input query. However, the matched rules exhibit varied relevance to the query with regards to their actual transformation. Specifically, the rules of the query rewrite engine can be classified into two types of transformations: the normalization rule (e.g., “SUB_QUERY_TO_JOIN”), which nearly always reduces query cost; and the exploration rule (e.g., “AGGREGATE_JOIN_TRANSPOSE”), which transforms the query but does not consistently result in cost reduction (Taft et al., 2020). We classify the rules based on expert experience, and assign relevance scores to the matched rules based on their transformation types. Initially, every rule has a score of $\infty$- ∞. Then, the matched normalization rules are assigned +$\infty$+ ∞ as closely relevant, and the matched exploration rules are assigned 00 as weakly relevant.
$(ii)$ Besides the directly relevant rules, we also detect indirectly relevant rules using the retrieved Q&As. Specifically, for each pair of rule and Q&A, we leverage LLM to evaluate whether the rule can be applied in the context of the Q&A. If applicable, the rule’s score is incremented by the Q&A similarity score to the input query. Note that if the rule has a score of $\infty$- ∞, we first initialize it as 00 before increment. Enumerating all the pairs of rules and the retrieved Q&As, we obtain the final score of each rule $r$. Note that the relevance can be performed by LLM offline for each possible pair of rule and Q&A, thus not affecting the algorithm’s efficiency.
$(iii)$ Given the rules with relevance scores, we first filter out those with $\infty$- ∞ as irrelevant. Then, we order the remaining rules by prioritizing the rules with higher scores, which comprehensively reflect their direct and indirect relevance to the input query.
Step 2: Recipe-based Rule Selection and Ordering. Derived recipes offer more detailed and comprehensive guidance for rewriting the input query when compared to the basic rule specifications and Q&As. Therefore, we refine the initially arranged rule sequence to align it more closely with these recipes. This refinement process is divided into three sub-steps.
$(i)$ We first refine rule selection with recipes. A straightforward method is to assess the relevance between each rule and each recipe, excluding those rules deemed irrelevant. However, this method faces two significant drawbacks. Firstly, it overlooks rule dependencies, consequently excluding indirectly utilized rules that become relevant only after another rule is applied. Secondly, this method incurs high costs, especially since it lies on the critical path for online query rewrite. To address these limitations, we propose a batch filtering method. Specifically, we first feed the rewrite recipe and a batch of rules to LLM, and require LLM to select the most appropriate rules aligned with the rewrite recipe, using the prompt $p_{select\_rule}=$“Given the input query and the rewrite rules, you should evaluate whether the rules can be applied to rewrite the query in the context of the recipe.” Next, we feed the previously selected rules along with the next batch of rules into LLM for further selection. We continue this process iteratively until all rules have been evaluated. In this way, we not only accelerate the LLM selection process, but also consider the rule dependency within the selected rules.
$(ii)$ Next, we methodically refine the rule order according to rewrite recipe. First, taking into account the inter-dependence of the selected rules, we categorize them into groups wherein each group pertains to the same SQL operator (e.g., join). We then instruct LLM to arrange them following the rewrite recipe, emphasizing the alignment of closely related rules. Second, leveraging the grouped rule ordering as a basis, we instruct LLM to refine the overall rule ordering to optimally align with the rewrite recipe.
$(iii)$ We input the refined rule sequence to the query rewrite engine, and rewrite the input query with the rules.
Step 3: Rewrite Reflection. With regards to the unstable performance of LLM in complex tasks like rule arrangement (Wei et al., 2022; Yao et al., 2024), we reflect the rewrite process to determine whether to finalize or refine the rewrite as discussed in Section 3.
Table 3. Comparison of Query Latency.
Query Latency (s) TPC-H 10x DSB 10x Calcite (uni)
Average Median p90 Average Median p90 Average Median p90
Origin 104.86 10.60 300.00 37.76 5.28 300.00 109.73 56.35 300.00
LearnedRewrite
67.12
($\downarrow$36.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
33.67
($\downarrow$10.8%)
4.67
($\downarrow$11.6%)
57.60
($\downarrow$80.8%)
12.63
($\downarrow$88.5%)
5.24
($\downarrow$90.7%)
48.47
($\downarrow$83.8%)
GPT-3.5
85.98
($\downarrow$18.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
37.75
($\downarrow$0.0%)
5.36
($\uparrow$1.5%)
300.00
($\downarrow$0.0%)
55.41
($\downarrow$49.5%)
22.74
($\downarrow$59.6%)
230.99
($\downarrow$23.0%)
GPT-4
67.10
($\downarrow$36.0%)
10.60
($\downarrow$0.0%)
300.00
($\downarrow$0.0%)
37.77
($\uparrow$0.0%)
4.92
($\downarrow$6.8%)
300.00
($\downarrow$0.0%)
60.86
($\downarrow$44.5%)
20.06
($\downarrow$64.4%)
300.00
($\downarrow$0.0%)
R-Bot (GPT-3.5)
56.24
($\downarrow$46.4%)
10.60
($\uparrow$0.0%)
300.00
($\downarrow$0.0%)
26.15
($\downarrow$30.7%)
4.61
($\downarrow$12.7%)
35.25
($\downarrow$88.2%)
37.71
($\downarrow$65.6%)
8.37
($\downarrow$85.1%)
65.67
($\downarrow$78.1%)
R-Bot (GPT-4)
57.60
($\downarrow$45.1%)
10.37
($\downarrow$2.2%)
300.00
($\downarrow$0.0%)
25.35
($\downarrow$32.9%)
4.58
($\downarrow$13.2%)
17.17
($\downarrow$94.3%)
12.45
($\downarrow$88.6%)
5.04
($\downarrow$91.0%)
48.30
($\downarrow$83.9%)
7. Experiments
We conduct extensive experiments to evaluate our system R-Bot and compare with state-of-the-art query rewrite systems.
7.1. Experiment Setting
We implement our system R-Bot using the rules in an open-sourced query engine Apache Calcite (Begoli et al., 2018). We execute the SQL queries in PostgreSQL v14. The server is a machine with 125.6 GB RAM and 3.1GHz CPU.
Datasets. To verify the effectiveness of R-Bot on different scenarios, we conduct experiments on three types of datasets. $(i)$ TPC-H is a standard OLAP benchmark, which contains 62 columns and 44 queries. We separately test R-Bot on different data sizes, i.e., TPC-H 10x ($\sim$10G) and TPC-H 50x ($\sim$50G). $(ii)$ DSB is a more complex OLAP benchmark (Ding et al., 2021), which contains 429 columns and 76 queries. We also test R-Bot on datasets DSB 10x ($\sim$10G) and DSB 50x ($\sim$50G). $(iii)$ Calcite is constructed from Apache Calcite’s rewrite rule test suites (cal, 2024), which is widely used in evaluating query rewrite capabilities (Wang et al., 2022; Dong et al., 2023; Bai et al., 2023). In our experiment, we select 44 representative queries with great rewrite potentials on a schema of 43 columns.  [^1] We chose queries from the Calcite test suites that showed the top 11% biggest performance improvements when rewritten by the Calcite query rewriter. Additionally, we evaluate R-Bot with a 10G data size across distinct data distributions, namely Calcite (uni) for uniform distribution and Calcite (zipf) for Zipfian distribution.
LLMs. We use the state-of-the-art LLM gpt-4o, and alternatively cheaper gpt-3.5-turbo-0125 (ope, 2024). We use the default temperature of 0.1, in balance of reproduction and LLM performance (lla, 2024).
Evaluation Metrics. We evaluate R-Bot using two metrics. $(i)$ Query Latency, which measures the execution time of the rewritten query. $(ii)$ Overall Latency, which captures the total time spent on both rewriting and executing the query. For each query, we conduct five executions and calculate the average, excluding the highest and lowest values. For each metric, we evaluate performance based on the average, median, and 90th percentile (p90).
Evaluated Methods. The evaluated methods include: $(i)$ R-Bot (GPT-4) is the version of R-Bot using gpt-4o. $(ii)$ R-Bot (GPT-3.5) is the version of R-Bot using gpt-3.5-turbo-0125. $(iii)$ LearnedRewrite employs Monte Carlo Tree Search (MCTS) to explore the optimal sequence of rewrite rules (Zhou et al., 2021). It estimates the performance improvements of the rewritten queries by utilizing database query costs, which guides the search process.  [^2] The origin LearnedRewrite learns neural networks from historical queries to estimate the cost reduction of applying rewrite rules. However, as our experiments are conducted in a practical setting without training examples, we replace the learning-based cost model with database query costs, which is more accurate while incurring longer rewrite latency. $(iv)$ GPT-4 model without using techniques in R-Bot, which inputs the SQL query and the rewrite rules, and outputs the arranged rule sequence. $(v)$ GPT-3.5 model similarly without using techniques in R-Bot.
7.2. Performance Comparison
We compare R-Bot with two types of baselines, including learning-based methods (LearnedRewrite), and origin LLMs (GPT-4, GPT-3.5).
Query Latency Reduction. Table 3 shows the query latency of rewritten queries. R-Bot outperforms the other query rewrite methods across all the datasets and metrics. The reasons are two-fold.
First, R-Bot can judiciously retrieve relevant rule specifications and Q&As, and use them to comprehensively figure out the rewrite rules pertinent to the input query. For instance, R-Bot can utilize the rewrite evidences to decide to apply the exploration rule “EXPAND_DISTINCT_AGGREGATES_TO_JOIN” weakly related to the input query, which results in a query rewrite of 5.6x acceleration. However, the other methods do not figure out this critical rule without guidance of rewrite evidences. Besides, we find that GPT-4 and GPT-3.5 tend to select only a small number of rules (e.g., 1), which is often sub-optimal. That is because, they lack of rewrite evidences, and thus can hardly uncover the intricate correlations between the rewrite rules and the input query.
Second, R-Bot adopts a step-by-step LLM rewrite method, which can leverage rewrite evidences to understand the interrelations among the rewrite rules during rule arrangement. For instance, R-Bot can accurately select and arrange multiple rewrite rules (e.g., 6) from analysis of retrieved evidences, so that they can co-operate together to rewrite the input query to an execution-efficient form (e.g., 4.6x accelerated). On the contrary, since LearnedRewrite relies on blind exhaustive search, it can hardly find the optimal rule sequence among tremendous possibilities (e.g., only 4.3x accelerated).
Figure 5. Comparison of Overall Latency on TPC-H 50x.
Figure 6. Comparison of Overall Latency on DSB 50x.
Overall Latency Reduction. We also evaluate the overall latency of the methods on larger-scale datasets TPC-H 50x and DSB 50x. Figure 5 and Figure 6 demonstrate the results. We find that R-Bot achieves the best overall latency on the two datasets. The reasons are three-fold. First, R-Bot can still optimize the query latency by finding better plans, thus outperforming other methods. Second, R-Bot takes some time (e.g., average around 1 min) to retrieve beneficial evidences and iteratively instruct LLM to select and arrange rewrite rules. Thus it can steadily identify the critical rewrite rules (e.g., “FILTER_SUB_QUERY_TO_JOIN”) with much higher query latency reduction. Instead, other methods do not spend efforts understanding rewrite evidences, which often causes sub-optimal rewritten queries. Third, compared with results of TPC-H 10x and DSB 10x (see Table 3), we find that R-Bot remains similar rewrite latency but demonstrates higher query latency reduction proportional to data scale, which both results in superior overall latency.
Table 4. Comparison of Query Improvement Ratio.
Improve (#Queries) TPC-H 10x DSB 10x Calcite (uni)
LearnedRewrite
6/44
(13.6%)
9/76
(11.8%)
36/44
(81.8%)
GPT-3.5
3/44
(6.8%)
4/76
(5.3%)
16/44
(36.4%)
GPT-4
6/44
(13.6%)
4/76
(5.3%)
21/44
(47.7%)
R-Bot (GPT-3.5)
21/44
(47.7%)
16/76
(21.0%)
31/44
(70.4%)
R-Bot (GPT-4)
17/44
(38.6%)
19/76
(25.0%)
39/44
(88.6%)
Query Improvement Ratio. Table 4 shows the query-level latency reduction ratio by different query rewrite methods, with R-Bot ranging from about 21.0% to 88.6%. We find that R-Bot still outperforms the other methods on the three datasets. That is because R-Bot can discover targeted beneficial rule arrangements using relevant evidences, while the other methods have difficulty specifying the optimal arrangement among tremendous search space.
Table 5. Comparison of Query Latency on Calcite (zipf).
Query Latency (s) Calcite (zipf)
Average Median p90
Origin 106.31 37.91 300.00
LearnedRewrite
15.94
($\downarrow$85.0%)
4.96
($\downarrow$86.9%)
34.96
($\downarrow$88.3%)
GPT-3.5
58.33
($\downarrow$45.1%)
20.04
($\downarrow$47.1%)
300.00
($\downarrow$0.0%)
GPT-4
61.80
($\uparrow$41.9%)
14.15
($\downarrow$62.7%)
300.00
($\uparrow$0.0%)
R-Bot (GPT-3.5)
32.44
($\downarrow$69.5%)
6.58
($\downarrow$82.6%)
57.40
($\downarrow$80.9%)
R-Bot (GPT-4)
7.56
($\downarrow$92.9%)
4.96
($\downarrow$86.9%)
18.08
($\downarrow$94.0%)
Robustness. Table 5 demonstrates the robustness of query rewrite methods on the Calcite dataset, which transfers from uniform distribution to zipf distribution. We have two observations. First, compared with Calcite (uni) (see Table 3), the performance of R-Bot remains consistent. R-Bot still achieves the lowest average, median, and p90 of query latency among the methods. That is because the rewrite evidences can cover different query rewrite scenarios, and R-Bot can adaptively decide among potential rule arrangements with feedbacks of database query cost. Second, we find that LearnedRewrite achieves comparative median query latency with R-Bot on Calcite (zipf) dataset. That is because the queries in the Calcite dataset often have great rewrite potentials of applying the rules, where LearnedRewrite can easily find beneficial rule orders even with blind explorations.
Evaluation Across Various LLMs. As shown in Table 3, we find that R-Bot (GPT-3.5) achieves comparative performance with R-Bot (GPT-4). That is because the algorithm design of R-Bot always decomposes complex tasks into simpler stages, both in evidence retrieval and step-by-step LLM rewrite, where each stage can be manageable with inferior LLM (e.g., GPT-3.5).
7.3. Ablation Study
Table 6. Ablation Study of Rewrite Evidence.
Query Latency (s) Calcite (uni)
Average Median p90
Naive RAG 39.13 14.88 98.15
GPT-4 60.86 20.06 300.00
Figure 7. Distribution of Rewrite Evidences with Different SQL Operations.
Table 7. Ablation Study of Structure-Semantics Retrieval.
Query Latency (s) Calcite (uni)
Average Median p90
Structure-Semantics Retrieval 12.45 5.04 48.30
Structure-Only Retrieval 31.96 5.30 56.40
Semantics-Only Retrieval 39.45 8.37 65.67
Naive RAG 39.13 14.88 98.15
Figure 8. Impact of Retrieval Top-k on Calcite (uni).
Figure 9. Impact of Rewrite Reflection on Calcite (uni).
Table 8. Ablation Study of Step-by-Step LLM Rewrite.
Query Latency (s) Calcite (uni)
Average Median p90
Step-by-Step LLM Rewrite 12.45 5.04 48.30
One-Step LLM Rewrite 36.47 13.83 69.43
7.3.1. Rewrite Evidence Preparation.
We evaluate the essence of rewrite evidence by comparing GPT-4 with naive RAG. As shown in Table 6, naive RAG equipped with evidences can outperform GPT-4 across all the metrics, because naive RAG can retrieve rewrite evidences that are somewhat relevant to the input query, which provides beneficial guidance in rewrite rule selection and ordering.
Besides, we also investigate the coverage of prepared rewrite evidences in Figure 7, where 67 rewrite rule specifications are extracted (30 from 119 chapters of documents, 37 from rule codes), and 2091 rewrite Q&As are filtered out from millions of website Q&As. We find that the rule specifications and Q&As both involve a wide range SQL operations (e.g., joins, sub-queries), and can thus provide supportive evidence for various query rewrite scenarios.
7.3.2. Hybrid Structure-Semantics Evidence Retrieval.
To evaluate the effectiveness of hybrid structure-semantics Q&A retrieval, we evaluate R-Bot separately with semantics-only and structure-only retrieval. As indicated in Table 7 , both methods exhibit a decline in performance compared to R-Bot. This reduction is due to their inability to identify Q&As beneficial for query rewrite from a holistic perspective, as they either overlook query semantics or structure information. For example, compared with R-Bot, they may miss key evidence, leading to the omission of critical rules like “EXPAND_DISTINCT_AGGREGATES” in the rule arrangement.
Retrieval Top-k. We also evaluate the impact of retrieval top-$k$, which decides the number of retrieved Q&As. As shown in Figure 9, we have two observations with varying values of $k$. First, the rewrite latency increases as $k$ grows. This is because R-Bot tends to generate more recipes with a greater number of retrieved Q&As. Consequently, as the recipes compose longer LLM contexts, each invocation of LLM requires additional time to process the extended context. Second, as $k$ increases, query latency generally decreases between 1 and 10, but increases between 10 and 20. The reasons are two-fold. $(i)$ When $k$ is small, the additional retrieved Q&As are more likely to cover essential evidence, allowing R-Bot to select critical rules (e.g., “AGGREGATE_JOIN_TRANSPOSE”) to facilitate query rewrite. $(ii)$ For large $k$, it is likely to already cover the essential evidence in the retrieved Q&As, making further increase unnecessary as it may not provide additional relevant information. Besides, increasing $k$ too much may introduce noisy contexts, which may impair performance of LLM in query rewrite. Therefore, we set $k=10$ in the other experiments.
7.3.3. Step-by-Step LLM Rewrite.
We compare our step-by-step LLM rewrite with one-step LLM rewrite that directly instructs LLM to arrange the rules in one step using retrieved rule specifications and Q&As, and evaluate the performance. As shown in Table 8, we find that one-step LLM rewrite often selects a smaller number of rules, which results in sub-optimal rule arrangements. That is because, given the extremely long context of retrieved evidences and available rewrite rules, LLM tends to produce short-cut solutions overlooking many critical details, and thus only selects the rules explicitly related to the input query. Instead, R-Bot employs a step-by-step LLM rewrite for reasoning, enabling the selection of higher-quality rules.
7.3.4. Rewrite Reflection.
We also evaluate the impact of rewrite reflection in R-Bot. As shown in Figure 9, we have two observations. First, the rewrite latency increases due to the additional overhead introduced by rewrite reflection and LLM self-correction. Second, the query latency decreases with further rewrite reflection, leading to an overall latency reduction. That is because rewrite reflection helps mitigate the hallucination issue in LLMs by exploring multiple rewrite rule arrangements to find the most beneficial query rewrite.
8. Related Work
Query Rewrite. Existing query rewrite methods mainly adopt two paradigms. (1) Heuristic-based methods (mys, 2024; pos, 2024; Graefe and McKenna, 1993; Graefe, 1995; Begoli et al., 2018). Some methods (e.g., PostgreSQL (pos, 2024)) apply rewrite rules in a fixed order, which often overlook better orders in different scenarios. Besides, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to explore different rule orders with the aid of heuristic acceleration. However, by neglecting inter-dependencies among the rules, these methods perform a blind search, often failing to identify the optimal orders within reasonable time. (2) Learning-based methods. LearnedRewrite (Zhou et al., 2021) decides the rule order using Monte Carlo Tree Search guided by learned cost models. However, the learned model cannot be transferred to unseen database schema without additional retraining, which is often impractical. Hence, it calls for a query rewrite system capable of reliably identifying an appropriate sequence of rewrite rules to optimize query rewrite.
Automatic Rewrite Rule Discovery. Given the tedious effort required to curate a large number of high-quality rewrite rules, some methods have been recently proposed to automate the process of rule discovery (Wang et al., 2022; Dong et al., 2023; Ding et al., 2023). WeTune (Wang et al., 2022) first represents the condition and transformation of rewrite rule with operators (e.g., projection) and symbols (e.g., relation, attribute). Then it enumerates potential rewrite rules within a constrained set of transformations (e.g., involving up to four operators) and conditions (combinations of seven basic types), verifying rule equivalence using SMT-solvers (Zhou et al., 2022; Chu et al., 2017). Besides, SlabCity (Dong et al., 2023) discovers possible query rewrites for a specific query by recursively applying various transformations (e.g., expanding SQL conditions) to the original query, and preserves the equivalent rewrites. Complementary to the above works, R-Bot retrieves highly relevant evidences to instruct LLM, effectively arranges newly discovered rules, and achieves superior adaptability and robustness in rule selection and ordering.
LLM for Database. Recent advances in LLMs have demonstrated superiority in natural language understanding and programming, which offers opportunities to leverage vast amounts of database forums to enhance database performance (Achiam et al., 2023; Liu et al., 2024b; Fan et al., 2024; Lao et al., 2024; Zhou et al., 2024a, b). For instance, GPTuner (Lao et al., 2024) proposes to enhance database knob tuning using LLM by leveraging domain knowledge to identify important knobs and coarsely initialize their values for subsequent refinement. Besides, D-Bot (Zhou et al., 2024a) proposes an LLM-based database diagnosis system, which can retrieve relevant knowledge chunks and tools, and uses them to accurately identify typical root causes. Unlike these approaches, R-Bot is a novel query rewrite system powered by LLM.
9. Conclusion
In this paper, we proposed an LLM-based query rewrite system. First, we prepared rewrite evidences from diverse sources, including rewrite rule specifications and rewrite Q&As. Next, we proposed a hybrid structure-semantics retrieval method to retrieve relevant rewrite evidences, based on which we generated rewrite recipes to instruct LLM for query rewrite. Then, we proposed a step-by-step LLM method, which iteratively utilized the retrieved Q&As and rewrite recipes to select and arrange rewrite rules with self-reflection. Experimental results demonstrated that R-Bot achieved remarkable improvements over existing query rewrite methods.